Consider two stocks. Stock A has a standard deviation of 46% and stock B has a standard deviation of 52%. The stocks have a correlation of 0.78. You plan to invest $7,607 into stock A and $6,081 into stock B. What is the standard deviation of your two stock portfolio? (round weights to 3 decimal places and final answer to 2 decimal places).
Answer #1

weight of stock A w1= 7607/(7607+6081)

= 0.556

weight of stock B w2=0.444

standard deviation of portfolio = sqrt(w1^2 * s1^2 + w2^2 * s2^2 + 2 * w1 * w2 * s1 * s2 * corr)

= sqrt(0.556^2 * 0.46^2 + 0.444^2 * 0.52^2 + 2*0.556 * 0.444 *0.46 * 0.52 * 0.78)

= 45.92%